# Determining Magnetic and Electric Field Generations in Laser-Driven Coil Targets

### v1

@@ 1 Introduction
In this study, we conduct a systematic investigation of proton radiograph patterns produced by a laser-driven coil target, focusing on a case where the magnetic field originates from a coil current and the electric field arises from a uniform charge distribution along the coil. Using ray tracing simulations, we analyze proton deflection patterns along two orthogonal axes—face-on and side-on—and examine the respective contributions of magnetic fields, electric fields, and their combined effects. By identifying and characterizing key features in the deflection patterns for each field configuration, we develop an approach to estimate the magnetic field contribution from proton radiographs. We then apply this methodology to our experimental data and find strong consistency between our analysis and measured results[27]. Our study provides insights into how proton radiographs can be interpreted in laser-driven coil targets, contributing to a more refined understanding of field characterization in high-energy-density physics experiments.
@@ 3.2 Side on proton radiographs
where is a dimensionless coefficient and can be determined from simulations. Figure 5 shows the central grid rotation angle as a function of coil current for 10, 25 and 40 MeV. The scattered points represent simulation data, while the lines correspond to fits based on Equation 7. The coefficient is obtained by fitting the 25 MeV data and is used consistently for all three curves. The results demonstrate that Equation 7 accurately describes the relationship between rotation angle and current. In practical applications, the current is estimated from the measured rotation angle. Rearranging the equation, the current for the coil can be estimated as:
For targets with different shapes, the coefficients in Equations 7 and 8 may vary. They can be determined using similar techniques described here, adapted to the specific geometry and magnetic field configurations.
@@ 3.3 Determining the electromagnetic fields
Table 1 presents a summary of the proton radiograph features for different cases in face-on and side-on proton radiography configurations. From the features, one can qualitatively determine the dominant field contribution in an experimental proton radiograph.
@@ 3.3.3 Application to experiments
Using the side-on grid rotation method, we successfully measured a 120 kA current flowing through a double-coil target driven by a short-pulse laser [12, 27]. This current generates a 200 T magnetic field at the center of one circular arc. When we overlay the simulation results for the 120 kA current onto the experimental radiograph, we observe excellent agreement in both grid rotation and leg widths. To estimate the upper-bound charge density and electric field, we included an additional electric field in the simulation. The resulting upper-bound charge density is , corresponding to an upper-bound electric field strength of approximately V/m at a distance of 0.3 mm away from the coil. Additionally, this platform for generating strong magnetic fields using short-pulse lasers is also applied to magnetic reconnection studies. Strong particle acceleration is observed under this intense magnetic field. Further details will be published in another paper.
@@ 4 Conclusions
In summary, we systematically investigated the features of proton radiographs under varying electromagnetic field conditions using face-on and side-on configurations. By analyzing the distinct proton deflection patterns in these configurations, we identified key features attributable to magnetic and electric fields, as summarized in Table 1. Based on these findings, we propose methods to isolate and quantify the magnetic field contribution using the side-on central grid rotation and the face-on bubble and leg width analysis. The side-on central grid rotation method is particularly preferable, as the rotation is uniquely sensitive to the magnetic field. The electric field contribution can then be determined by comparing leg width differences between combined-field and magnetic-field-only cases. We applied the side-on central grid rotation method to evaluate the electromagnetic fields generated by a U-shaped double-coil target, demonstrating its effectiveness in enhancing the diagnostic capabilities of proton radiography for accurate and reliable field measurements in high-energy-density experiments. Moreover, the detailed feature analysis presented in this work provides a better understanding of the interplay between electric and magnetic field contributions in proton radiographs, facilitating the optimization of laser-driven capacitor coil designs for advanced applications. Looking ahead, this proton radiograph analysis could be extended to more complex field geometries, including those with non-uniform, dynamic, or time-varying current and charge distributions.

### v2

# Diagnosing electric and magnetic fields in laser-driven coil targets

@@ 1 Introduction
Previous studies have utilized face-on void (or bubble) size in proton radiographs to infer coil current and the associated magnetic field strength [5, 25, 21]. Since such voids can be produced either by a positively charged wire (via electric fields) or by current flowing through the wire (via magnetic fields), a single void size from one radiograph alone cannot unambiguously distinguish between the two.
Because electric and magnetic deflections exhibit different scaling with proton energy, for magnetic fields, the void size scales as [5], whereas for electric fields it scales as [22]. A dual-energy proton deflection method has been proposed to differentiate the field types [26]. While promising, this approach relies on the availability of a broad proton energy spectrum. The combination of high energy and wide energy spread—may not be readily achievable at all experimental facilities[25].
An alternative is to reverse the proton beam direction in a second shot, flipping the magnetic deflection while keeping the electric deflection unchanged. Vlachos et al. [24] demonstrated this for a magnetic-dominated case. This method requires two separate shots for a single measurement, and its reliability depends on maintaining low shot-to-shot variability.
Side-on grid rotation analysis offers another approach to infer coil current and magnetic field strength in capacitor-coil targets [20, 21, 22, 25]. Since the induced rotation angle depends solely on the magnetic field, this technique provides a direct and unambiguous measurement of magnetic field strength. As the proton beam is aligned with the main magnetic field direction in this configuration, the resulting deflections are smaller, leading to reduced field sensitivity.
In this study, we conduct a systematic investigation of proton radiograph patterns produced by a laser-driven coil target, focusing on a case where the magnetic field originates from a coil current and the electric field arises from a uniform charge distribution along the coil. Using ray tracing simulations, we analyze proton deflection patterns along two orthogonal axes—face-on and side-on—and examine the respective contributions of magnetic fields, electric fields, and their combined effects. By identifying and characterizing key features in the deflection patterns for each field configuration, we develop an approach to estimate the magnetic field and electric field contributions from proton radiographs. We then apply this methodology to our experimental data and find strong consistency between our analysis and measured results[27]. Our study provides insights into how proton radiographs can be interpreted in laser-driven coil targets, contributing to a more refined understanding of field characterization in high-energy-density physics experiments.
@@ 3.2 Side on proton radiographs
where is a dimensionless coefficient and can be determined from simulations. Figure 5 shows the central grid rotation angle as a function of coil current for 10, 25 and 40 MeV. The rotation angle is defined over the central region of the coil—from to , centered at —where the rotation angle remains relatively constant. The scattered points represent simulation data, while the lines correspond to fits based on Equation 7. The coefficient is obtained by fitting the 25 MeV data and is used consistently for all three curves. The results demonstrate that Equation 7 accurately describes the relationship between rotation angle and current. In practical applications, the current is estimated from the measured rotation angle. Rearranging the equation, the current for the coil can be estimated as:
For coil targets with different geometries, the numerical coefficients in Eqs. 7 and 8 may change. These coefficients can be recalculated using the same approach described here, tailored to the specific geometry and magnetic field profile.
@@ 3.3 Determining the electromagnetic fields
Figure 6 presents a summary table of the proton radiograph features for different cases in face-on and side-on proton radiography configurations. From the features, one can qualitatively determine the dominant field contribution in an experimental proton radiograph.
@@ 3.3.1 Side-On Grid Rotation Method
To apply the grid rotation method, a sufficiently large rotation angle is required—typically at least 1 degree for reliable measurement based on our experiment. This threshold may vary across different experimental setups, as the minimum resolvable rotation angle depends on factors such as proton energy, coil geometry, and the relative positions of the proton source, coil, and detector. For the coil geometry discussed in this paper, the minimum measurable current under 1 degree rotation must satisfy where is the proton beam energy. For lower current scenarios, using a lower-energy proton beam is feasible.
@@ 3.3.3 Application to experiments
Our approach introduces two complementary diagnostic methods that extract both magnetic and electric field components from a single radiograph, in either side-on or face-on geometry. Each method leverages two independent and physically distinct observables, forming a coupled system of equations that can be uniquely solved. This strategy eliminates the need for multiple shots or broad proton energy spectra, while providing quantitative access to both magnetic and electric field contributions in a single experiment.
Using the side-on grid rotation method, we successfully measured a 120 kA current flowing through a double-coil target driven by a short-pulse laser [12, 27]. This current generates a 200 T magnetic field at the center of one circular arc. When we overlay the simulation results for the 120 kA current onto the experimental radiograph, we observe excellent agreement in both grid rotation and leg widths. To estimate the upper-bound charge density and electric field, we included an additional electric field in the simulation. The resulting upper-bound charge density is , corresponding to an upper-bound electric field strength of approximately V/m at a distance of 0.3 mm away from the coil. Additionally, this platform for generating strong magnetic fields using short-pulse lasers is also applied to magnetic reconnection studies. Strong particle acceleration is observed under this intense magnetic field. Further details can be found in another paper[27].
3.3.4 Comparison of Electric and Magnetic Deflection Contributions
The characteristic field strengths scale as: , , where is the coil current, is the line charge density, and is the distance from the source. The ratio of electric to magnetic deflection for a beam with velocity orthogonal to the magnetic field direction can then be estimated as:
Here, in face-on configuration and in side-on direction.
This expression indicates that when , the magnetic deflection dominates; whereas if , the electric deflection becomes dominant. This estimate provides a practical method for determining the electric and magnetic contributions from line charge density and current.
@@ 4 Conclusions
In summary, we systematically investigated the features of proton radiographs under varying electromagnetic field conditions using face-on and side-on configurations. By analyzing the distinct proton deflection patterns in these configurations, we identified key features attributable to magnetic and electric fields, as summarized in Figure 6. Based on these findings, we propose methods to isolate and quantify the magnetic field contribution using the side-on central grid rotation and the face-on bubble and leg width analysis. The side-on central grid rotation method is particularly preferable, as the rotation is uniquely sensitive to the magnetic field. The electric field contribution can then be determined by comparing leg width differences between combined-field and magnetic-field-only cases. We applied the side-on central grid rotation method to evaluate the electromagnetic fields generated by a U-shaped double-coil target, demonstrating its effectiveness in enhancing the diagnostic capabilities of proton radiography for accurate and reliable field measurements in high-energy-density experiments. Moreover, the detailed feature analysis presented in this work provides a better understanding of the interplay between electric and magnetic field contributions in proton radiographs, facilitating the optimization of laser-driven capacitor coil designs for advanced applications. Looking ahead, this proton radiograph analysis could be extended to more complex field geometries, including those with non-uniform, dynamic, or time-varying current and charge distributions.